OpenStudy (anonymous):

The height of the tunnel at the center is 27 ft and the vertical clearance must be 9 ft at a point 24 ft from the center. Find an equation for the ellipse. Is this correct Center = (0,0) | Points = (0,27), (-24,9), (24,9) | Vertex = (0,27) (y/a)^2 + (x/b)^2 = (9/27)^2 + (24/b)^2 = 1 b^2 = 648 --- y^2/729 + x^2/648 = 1

OpenStudy (neer2890):

ellipse with centre at (0,0) and major axis =2a and minor axis =2b \[\frac{ x ^{2} }{ a ^{2} }+\frac{ y ^{2} }{ b ^{2} }=1\] 

OpenStudy (anonymous):

This is what I have now: 24^2/b^2 + 9^2/27^2 = 1 24^2/b^2 = 8/9 24^2/*b^2 = 8/9 8/9*b^2 = 576 b^2 = 648 --- x^2/648 + y^2/729

OpenStudy (mathmate):

Your answer is mathematically correct. The best way to solve this type of problems is to do what @neer2890 did: draw a diagram and put numbers in so you can visualize the problem and anticipate the solution graphically. This will help you visualize your next problem, and come up with a logical solution.  This looks like a wide tunnel which is rather expensive to construct. Double-check the question to see if the tunnel is 24 ft. high at 9 ft from the centre!